Suppose z is a function of x and y, and tany2+x2 = zxe6y. Solve for ∂z/∂x and ∂z/∂y.


1
Expert's answer
2022-05-13T05:19:06-0400
"\sqrt{y^2+x^2}=z^xe^{6y}"

"\dfrac{1}{2}\ln(y^2+x^2)=x\ln z+6y"

Differentiate both sides with respect to "x"


"\dfrac{2x}{2(y^2+x^2)}=\ln z+\dfrac{x}{z}(\dfrac{\partial z}{\partial x})"


"\dfrac{\partial z}{\partial x}=\dfrac{z}{y^2+x^2}-\dfrac{z\ln z}{x}"



"\dfrac{1}{2}\ln(y^2+x^2)=x\ln z+6y"

Differentiate both sides with respect to "y"


"\dfrac{2y}{2(y^2+x^2)}=\dfrac{x}{z}(\dfrac{\partial z}{\partial y})+6"

"\dfrac{\partial z}{\partial y}=\dfrac{yz}{x(y^2+x^2)}-\dfrac{6z}{x}"